OpenStudy (anonymous):

a 50 milliliter solution of acid and water contains 25% acid. how much water do you need to add in order to make a 10% acid solution?

OpenStudy (inkyvoyd):

Eh, I'm not good at this. I can tell you how I do it, but It's probaly not he most effiecient

OpenStudy (inkyvoyd):

50*0.25=0.1x

OpenStudy (inkyvoyd):

And that's assuming the 25% is by volume, not by weight

OpenStudy (inkyvoyd):

IF it's by weight, there is not enough info to do theproblem, so it's most likely by volume

OpenStudy (inkyvoyd):

Wait, 0.1 x is the total amount

OpenStudy (inkyvoyd):

50*0.25=0.1x then subtract 50 from x

OpenStudy (inkyvoyd):

@iHeartYou ?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you have 50 ml which is 25% acid, so you have \(.25\times 50=12.5\) ml of acid if you add say "x" amount of water you will have 50 + x ml of water and still have 12.5 ml of acid. you want this 12.5 ml of acid to be 10% of the total, so put \(.10(50+x)=12.5\) and solve for x

OpenStudy (anonymous):

you should get \(5+.1x=12.5\) the subtract 5 to get \(.1x=7.5\) divide by .1 to get \[x=7.5\div .1=75\div 1=75\] and so you need to add 75 ml of water

OpenStudy (anonymous):

you people are very smart